Project

General

Profile

09/17/2014

Attendance

Mat, Gavin, Pavan, Fernanda, Kyle, Michael

Far Detector

Diblock 4 is reinstalled and retrofit complete. Retrofitting Far Detector Diblock 5 Top.
Planned Network outage all day Tuesday 16th.
Promote the Detector Configuration to the checklist plots tab.
Which plots are looked at and which should be removed? Dedicate some time to identify plots, and which
plots do we use for diagnosis.
GoodRuns metric shows we are 99.8% efficient.

Improvements

Fernanda had several good ideas to promote better use of plots for shifter purposes. Volunteered to make some
slides that highlight the plots we use in the checklist and what are useful.
We will invite Rob/Jaroslav to watchdog meeting to talk them through the tools we have available and start
process of streamlining things.
Action Item: Mat will send email invite to Rob/Jaroslav.
Rob appears to have identified new Fermilab postdoc Keith Matera as the conduit between data quality/watchdog (re: outfitting meeting).

DCM Sync

DCM Sync plots. Jose will give an update on Friday.
Action Item: Follow up with Jose at least at Friday's meeting.

DDT

Spread in DDT plots seems odd. Why is it not more stable, especially for the calibration trigger?
Kyle is asking DDT group

Multi-point

September 11th started reading out running 4-sample multi-point (nanoslice V3) on FD.
Some Nearline plots showed different behavior related to reconstructed slices [ECL 49147], after Fernanda deployed better timing delay constants (DCM offsets) things got better. Are all changes related to multi-point or also due to software upgrades?
All processing moved to new tag. P2 is not using multi-point so this is purely using new software upgrades and there are no visible changes thus multi-point has caused the slight increase in some rates.
Best plot that highlights the change is standard deviation of hit times.

Hardware Watchlist for FD has rolled to new release. Plots are updating correctly, but debugging issue whereby the csv files are not updating correctly.
No new information since 09/10/14.
Action Item: Fernanda will fix the updating of the FD csvs.

Near Detector

Louise reported that ND had been running low gain since September 5th through ??th. Unclear from nearline how we would have spotted this.
Action Item: Ask Louise how she spotted this and how we are supposed to spot this going forward.


From outfitting meeting, the config was updated only today at noon so we will have to look at the plots between 5th- today (17th) and going forward to see any differences.

No reconstruction plots since they are only processed using numi triggers. Seems non-optimal going forward.
Action Item: What should we use to reco going forward? Will be addressed at Friday DQ meeting.

Trigger Server

Not updated since 12th.

NDOS

Action Item: Fernanda to make sure all NDOS links are removed. Follow up with Nitin.